OpenStudy (anonymous):

rationalize the denomiator: sqrt(4/x)

OpenStudy (inkyvoyd):

sqrt(4/x) =sqrt(4)/(sqrt(x) =[sqrt(x)sqrt(x)]/[sqrt(x)*sqrt(x)]

OpenStudy (diyadiya):

Multiply and divide denominator and numerator by \(\sqrt{x}\)

OpenStudy (anonymous):

what about this one \[4/ \sqrt[4]{3}\]

OpenStudy (inkyvoyd):

square both sides twice.

OpenStudy (anonymous):

could you show me

OpenStudy (diyadiya):

\[\frac{4}{\sqrt[4]{3}}\]\[\sqrt[4]{3}= (3)^ \frac{1}{4}\] it can be written as \[\frac{4}{(3)^ \frac{1}{4}}\] Multiply and divide by \( 3^\frac{1}{4}\) \[{4 \times 3^ \frac{1}{4} \over 3^ {\frac{1}{4}} \times 3^ \frac{1}{4}} \]

OpenStudy (diyadiya):

\[= {4 \times 3^ \frac{1}{4}} \over {3^\frac{1}{2} }\] Now again multiply and divide by \( 3^\frac{1}{2} \) \[\frac{ 4 \times 3^ \frac{1}{4} \times 3^ \frac{1}{2} }{ 3^ \frac{1}{2} \times 3^ \frac{1}{2}}=\frac{ 4 \times 3^ \frac{3}{4} }{3}\]